OpenStudy (anonymous):

Solve for x: 3(x+1)=-2(x-1)-4

OpenStudy (anonymous):

It can't be -9 it's not one of the answer choices :/

OpenStudy (anonymous):

A.1 b.-1 c.-5 d.-25

hero (hero):

3(x + 1) = -2(x - 1) - 4 3(x - 1 + 2) = -2(x - 1) - 4 3(x - 1) + 6 = -2(x - 1) - 4 3(x - 1) + 2(x - 1) = - 4 - 6 5(x - 1) = - 10 x - 1 = -10/5 @Vertik, finish it

OpenStudy (anonymous):

-1?

hero (hero):

Good job